OpenStudy (islandchick60):

Use the Zero Product Property to solve the equation. (x-9)(x-8)=0

OpenStudy (anonymous):

The "Zero Product Property" says: If (x−9)(x−8) = 0 then (x−9) = 0 or (x−8) = 0 So solve each of the equations for x.

OpenStudy (islandchick60):

so would it be x= -9 or -8?

OpenStudy (islandchick60):

@Zelda

OpenStudy (anonymous):

No. Your signs are wrong.

OpenStudy (islandchick60):

So x=9 or8

OpenStudy (islandchick60):

@Zelda can i ask you another question

OpenStudy (anonymous):

Dunno, can you?

OpenStudy (islandchick60):

Haha i guess so.

OpenStudy (islandchick60):

kk (4k+5)(k+7)=0 would that equal k=-5/4 or -7 ?

OpenStudy (anonymous):

Yes.

OpenStudy (islandchick60):

Yay! thx